### v1

# Lost in the Prompt Order: Revealing the Limitations of Causal Attention in Language Models

Abstract
Large language models exhibit surprising sensitivity to the structure of the prompt, but the mechanisms underlying this sensitivity remain poorly understood. In this work, we conduct an in-depth investigation on a striking case: in multiple-choice question answering, placing context before the questions and options (CQO) outperforms the reverse order (QOC) by over 14%p, consistently over a wide range of models and datasets. Through systematic architectural analysis, we identify causal attention as the core mechanism: in QOC prompts, the causal mask prevents option tokens from attending to context, creating an information bottleneck where context becomes invisible to options.
Lost in the Prompt Order:
Revealing the Limitations of Causal Attention in Language Models
Hyunjong Ok1,2 Jaeho Lee1 1POSTECH 2HJ AILAB hyunjong.ok@gmail.com, jaeho.lee@postech.ac.kr
1 Introduction
Large language models (LLMs) are highly sensitive to prompt structure. Even minor changes in surface forms—such as instruction phrasing, example placements, or reasoning elicitation—can lead to substantial differences in the prediction quality of the models wei2022chain; kojima2022large.
Despite the importance of this sensitivity for the practical reliability of LLMs, our current understanding remains largely descriptive. We know “what” LLMs are sensitive to, but not “why.” For instance, lu2022fantastically has shown that permuting the order of demonstrations in in-context learning (ICL) can dramatically affect accuracy, yet offers limited insight into what makes certain orders preferable. Likewise, recent studies on multiple-choice question answering (MCQA) report significant performance fluctuations under option permutations, but stop short of identifying the mechanisms underlying this phenomenon (pezeshkpour-hruschka-2024-large; zheng2024large).
In this work, we undertake an in-depth study of a less obvious but consequential form of prompt-sensitivity: the ordering of components in MCQA prompts. A typical MCQA prompt consists of three elements: a context passage (C), a question (Q), and a set of options (O). Intuitively, reordering these components should have little effect, as their semantic content remains unchanged. Contrary to this expectation, we find that placing the context before the questions and options (CQO) consistently and substantially outperforms the reverse ordering (QOC) across a wide range of setups (Figure˜1).111A similar phenomenon has also been noted by shaier2024not outside the MCQA context, which also does not demystify why such phenomenon happens.
To explain this phenomenon, we formulate three competing hypotheses and evaluate them through a series of carefully controlled experiments. By systematically validating (or invalidating) each hypothesis, we aim to uncover the underlying factors that drive context-order sensitivity in MCQA. Specifically, we consider the following hypotheses:
Hypothesis 1: Biased training data. CQO-style prompts may be more prevalent in training data, making QOC an unfamiliar format for the model.
Hypothesis 2: Failures in option recall. QOC structure makes it difficult for the model to recall the options located in the middle of the prompt, a.k.a. “lost-in-the-middle” (liu2024lost).
Hypothesis 3: Causal attention (winner). The causal attention structure in decoder-only transformers makes it impossible for the option tokens to attend to the context tokens, creating an information bottleneck (Figure˜2).
Our experiments—conducted on a wide range of datasets and models — support the third hypothesis, and rule out the other two hypotheses. Stepping further, we identify several factors that can affect the impact of causal masking, namely the context length and the option positions.
Finally, building on the causal-attention-based explanation, we design targeted interventions that can improve the performance of LLMs on QOC-structured prompts, or conversely, degrade the performance on CQO-structured prompts. These results provide additional evidence that causal attention is the driving mechanism of the sensitivity.
2 Experimental setup
Datasets.
We evaluate on four reading comprehension benchmarks that require context-based reasoning: LogiQA (liu2020logiqa); RACE-H/M (lai2017race); SciQ (welbl2017crowdsourcing). All tasks are in MCQA format with four options.
Models.
We conduct experiments on 21 decoder-only LLMs from four model families: LLaMA 3 (grattafiori2024llama), Qwen 2.5/3 (qwen2025qwen25technicalreport), and Gemma 2 (gemma_2024). Model sizes range from 0.5B to 9B parameters, including both base and instruction-tuned variants. For architecture comparison experiments (§3.3), we additionally test Flan-T5 (chung2024scaling) family (encoder-decoder) and BERT, RoBERTa, ALBERT (devlin2019bert) family (encoder-only).
Evaluation metrics.
We measure accuracy under two prompt orderings: CQO and QOC, and utilize the performance gap (), which quantifies context order sensitivity.
Other details.
Further details on datasets, models, prompting templates, and evaluation/scoring protocols are provided in Appendix B.
3 Hypotheses and analysis
As shown in Figure˜1, permuting the prompt order leads to a sharp performance drop across all benchmarks. To understand the cause, we formulate and test a set of hypotheses. All detailed experiments for each model and dataset are in Appendix˜C.
3.1 Hypothesis 1: Bias in training samples
Hypothesis. CQO is simply more frequent in the training data than QOC. Thus, in the QOC format, the model may fail to process an unfamiliar format.
Experiment. We test the training-distribution hypothesis in two ways. First, CQO-like prompts are more common in instruction data; this hypothesis predicts a larger CQO-QOC gap for instruction models. We therefore compare nine matched base-instruct pairs. Second, we use in-context learning to familiarize models with the QOC format, varying the number of demonstrations up to 5-shot.
Results. As shown in Figure˜3(a), the CQO-QOC gaps are nearly identical, suggesting the phenomenon is not driven by instruction tuning. Moreover, even with 5-shot demonstrations shown in Figure˜3(b), QOC accuracy improves by only 3.1% and remains far below CQO. Together, these results rule out training distribution as the primary cause.
3.2 Hypothesis 2: Failure to recall options
Hypothesis. LLMs often fail to recall the information located in the middle of the context (liu2024lost). Thus, in the QOC format, the model may fail to correctly recall the options, which are located between the question and the context.
Experiment. After presenting the prompt, we ask the LLMs to recall each option. Precisely, we measure the chance of an exact match for each option.
Results. As shown in Figure˜4, QOC achieves a similar, or even higher, recall accuracy than CQO. This indicates that the failure of option recall may not be the cause of accuracy drops in QOC.
3.3 Hypothesis 3: Causal attention
Hypothesis. Causal attention mask prevents option tokens from attending to context in QOC. In decoder-only, each token can only attend to preceding tokens. Thus, in QOC, where options appear before context, option representations are computed without information from the context.
Experiment 1: Architecture comparison. If causal masking is the root cause, models with bidirectional attention should exhibit no performance gap. We compare three architecture types: decoder-only (causal), encoder-decoder (bidirectional encoder), and encoder-only (bidirectional). For encoder-decoder models, we feed the entire prompt to the encoder and let the decoder generate the answer.
Results 1. As shown in Figure˜5(a), decoder-only models show a 14.72% gap, encoder-decoder models (Flan-T5) show 2.30%, and encoder-only models (BERT, RoBERTa, ALBERT) show a near-zero gap (0.02%). This pattern strongly implicates causal masking as the underlying mechanism.
Experiment 2: Context removal test. If the context is effectively inaccessible in QOC, then removing it entirely should not materially change performance. We therefore compare QOC against QO (Question-Options only), where the context is omitted.
Results 2. As shown in Figure˜5(b), QOC accuracy is nearly identical to QO, consistent with the model failing to use the context in QOC. This supports the hypothesis that the issue is not forgetting but restricted access induced by the causal mask.
Experiment 3: Attention analysis. We analyze how attention to option tokens evolves across layers. Also, we measure token contributions using GradientInput attribution (shrikumar2017learning), aggregating per-token scores over the context span and normalizing by total attribution.
Results 3. Option tokens receive exactly zero attention from context tokens in QOC by construction of the causal mask, confirming that the context-to-option pathway is blocked. Across layers, attention to options increases with depth in QOC but decreases in CQO, suggesting that QOC increasingly relies on the options themselves while CQO more effectively integrates contextual information (see Figure˜6(a)). Also, context attribution is 0.797 in CQO but only 0.335 in QOC (see Figure˜6(b)).
Why doesn’t the final attention fix it?
One may argue that the final answer token can still attend to the context in QOC. However, causal masking prevents the options themselves from being context-conditioned: all option hidden states are computed before any context is seen, so they cannot encode evidence–option alignment. Since later tokens cannot retroactively update earlier states (they can only read them), the model must do evidence–option comparison only at the final decoding step using context-blind option representations.
Modulating factors.
Two factors modulate the gap severity. First, longer contexts show larger gaps (Table˜1(a)): more context means more information becomes inaccessible under causal masking. Second, earlier answer positions suffer greater accuracy drops than later ones in QOC (Table˜1(b)), as later options are near the context.
| Short context | Long context | |||
| Dataset | LogiQA | SciQ | RACE-M | RACE-H |
| Avg. Length | 70 | 70 | 195 | 305 |
| Gap () | 6.2% | 7.3% | 24.8% | 20.8% |
| Position | A | B | C | D |
| Gap () | 22.4% | 20.5% | 19.2% | 9.9% |
4 Targeted interventions
We design targeted interventions that directly manipulate the option and context attention pathway (see Table˜2). We apply each intervention to all 21 decoder-only models across four datasets.
Attention pruning (degrading CQO).
To simulate the constraint imposed by causal masking in QOC, we block option-to-context attention in CQO. For each sample, we identify token spans for the context and the options, and set for all pairs where and . This prevents option tokens from attending to context tokens while leaving all other attention unchanged. CQO accuracy drops from 69.26% to 42.46% (), with consistent decreases across model families (Qwen: %, LLaMA: %, Gemma: %), indicating that option access to contextual evidence is critical.
Activation patching (improving QOC).
We restore context-aware option representations in QOC by patching option hidden states with those computed under CQO. For each sample, we run both templates and align corresponding option token positions by exact string matching. At layers in the middle-to-late half of the network (e.g., layers 12–23 for 24-layer models, normalized by depth), we replace with . We patch only option tokens (not context or question) to isolate the mechanism; patched states come from a different template on the same sample, and token alignment is verified by exact string match. This increases QOC accuracy by points on average, with larger gains for models exhibiting larger baseline gaps.
Option repetition (improving QOC).
As a simple prompt, we repeat the options after the context (QOCO). The repeated option tokens can attend to the context under the causal mask, requiring no internal model intervention. Repeating options improves QOC by points, partially closing the gap without modifying model internals.
| Methods | LogiQA | SciQ | RACE-M | RACE-H | Avg |
| CQO | 39.08 | 94.16 | 74.32 | 69.48 | 69.26 |
| + Attention pruning | 30.47 | 60.94 | 39.52 | 38.89 | 42.46 (-26.8) |
| QOC | 32.94 | 86.89 | 49.57 | 48.76 | 54.54 |
| + Activation patching | 35.18 | 87.64 | 62.22 | 56.94 | 60.49 (+6.0) |
| + QOCO (Repeat) | 35.62 | 91.80 | 63.55 | 59.65 | 62.76 (+8.2) |
5 Conclusion
We investigated why decoder-only LLMs exhibit a huge performance gap between CQO and QOC orderings in MCQA tasks. After ruling out training distribution bias and memory decay, we identified causal attention as the core mechanism: in QOC, the causal mask prevents options from attending to context, rendering it effectively inaccessible. This finding is supported by architecture comparisons (encoder models show no gap), attention analysis, and targeted interventions that successfully degrade CQO or improve QOC performance. Our work provides mechanistic insight into prompt sensitivity and offers practical guidance.
Limitations
While our study provides mechanistic insights into prompt order sensitivity, our evaluations are primarily focused on MCQA tasks with four-option formats. However, considering that the identified root cause—causal attention masking—is fundamental to decoder-only architectures, we believe our findings regarding the information bottleneck are likely applicable to broader QA formats where context placement varies. Additionally, although we evaluate models up to 9B parameters due to computational constraints, the consistent performance gap observed across 21 diverse models suggests that this behavior is intrinsic to the causal decoder design rather than a scale-dependent artifact. Our work establishes a rigorous empirical foundation for understanding prompt sensitivity.
Appendix A Related work
Prompt sensitivity.
A growing body of work has established that LLMs can be sensitive to prompt design choices, including input order and formatting. In the in-context learning (ICL) setting, lu2022fantastically show that reordering the same set of demonstrations can substantially change accuracy and propose strategies to mitigate order sensitivity, but do not explore the underlying internal mechanism. In MCQA, prior work reports large variation under option permutations and develops mitigation methods (e.g., pezeshkpour-hruschka-2024-large; zheng2024large), yet largely remains at the level of behavioral characterization. Prompt formatting studies further show that semantically preserving changes—punctuation, spacing, labeling, or layout—can induce sizable performance swings (e.g., sclar2024quantifying). More mechanistic MCQA analyses, such as wiegreffe2025answer, use activation patching-style interventions to localize where answer selection is determined and how it is amplified across layers. However, these lines of work do not directly address a block-order question that arises with long contexts: whether and how the accessibility of context information to option representations changes with prompt order. Relatedly, shaier2024not reports large order effects in reading comprehension, but stops short of providing a structural account of the failure mode. In contrast, we study a striking order effect in MCQA between CQO and QOC and identify an architectural mechanism: under causal attention, option tokens in QOC are structurally prevented from integrating information from the subsequent context. We then test this mechanism with targeted causal interventions that manipulate the option–context information pathway.
Mechanistic analysis tools.
To connect prompt-order sensitivity to internal computation, we draw on a set of mechanistic interpretability tools. First, we use attention analysis to characterize how models route information under different prompt orders; prior work has shown the utility of interpreting attention weights as explanations (clark2019does; abnar2020quantifying). We compute attention statistics for each prompt order, use them as descriptive diagnostics, and also perform gradient-based analyses. We employ gradient-based input attribution as an auxiliary diagnostic: GradientInput and related attribution methods quantify how sensitive the prediction is to perturbations of each input token representation, and have been widely used as simple, interpretable feature-importance baselines (shrikumar2017learning; ding2021evaluating; poche-etal-2025-consim). We also use causal interventions on internal pathways. In particular, we (i) ablate the option-to-context attention pathway by masking attention edges between token groups (a targeted path/edge ablation), and (ii) apply activation patching (causal tracing/interchange interventions), which swaps hidden states from a “clean” run into a “corrupted” run to test whether a specific representation is causally responsible for the downstream behavior (vig2020investigating; geiger2021causal; meng2022locating). We adapt these tools to the prompt-order setting by isolating the option–context information pathway: attention ablation tests necessity, and activation patching tests sufficiency of context-conditioned option representations for closing the CQO-QOC gap.
Long-context failure modes.
Prior work shows that LLMs can underutilize long contexts and are sensitive to where relevant evidence appears (e.g., lost-in-the-middle; liu2024lost; li2024loogle; an2024make). In our setting, we use long contexts to test a specific hypothesis: QOC may underperform because the model forgets the options after reading a long context. Our option-recall results reject this explanation—QOC retains options at least as well as CQO—suggesting that the CQO–QOC gap is not primarily a long-context memory failure, but a distinct prompt-order effect.
Appendix B Detailed information
To demonstrate the robustness of our findings, we conduct experiments with various models and datasets. Details are below.
B.1 Model information
| Model | Hugging Face ID |
| Qwen 2.5 Family | |
| Qwen2.5-0.5B | Qwen/Qwen2.5-0.5B |
| Qwen2.5-0.5B-Instruct | Qwen/Qwen2.5-0.5B-Instruct |
| Qwen2.5-1.5B | Qwen/Qwen2.5-1.5B |
| Qwen2.5-1.5B-Instruct | Qwen/Qwen2.5-1.5B-Instruct |
| Qwen2.5-3B-Instruct | Qwen/Qwen2.5-3B-Instruct |
| Qwen2.5-7B | Qwen/Qwen2.5-7B |
| Qwen2.5-7B-Instruct | Qwen/Qwen2.5-7B-Instruct |
| Qwen 3 Family | |
| Qwen3-0.6B | Qwen/Qwen3-0.6B |
| Qwen3-1.7B | Qwen/Qwen3-1.7B |
| Qwen3-4B | Qwen/Qwen3-4B |
| LLaMA 3 Family | |
| Llama-3.2-1B | meta-llama/Llama-3.2-1B |
| Llama-3.2-1B-Instruct | meta-llama/Llama-3.2-1B-Instruct |
| Llama-3.2-3B | meta-llama/Llama-3.2-3B |
| Llama-3.2-3B-Instruct | meta-llama/Llama-3.2-3B-Instruct |
| Llama-3.1-8B | meta-llama/Llama-3.1-8B |
| Llama-3.1-8B-Instruct | meta-llama/Llama-3.1-8B-Instruct |
| Gemma 2 Family | |
| gemma-2-2b | google/gemma-2-2b |
| gemma-2-2b-it | google/gemma-2-2b-it |
| gemma-2-9b | google/gemma-2-9b |
| gemma-2-9b-it | google/gemma-2-9b-it |
| Encoder-Decoder (Architecture Comparison) | |
| Flan-T5-small | google/flan-t5-small |
| Flan-T5-base | google/flan-t5-base |
| Flan-T5-large | google/flan-t5-large |
| Flan-T5-xl | google/flan-t5-xl |
| Encoder-only (Architecture Comparison) | |
| BERT-base | google-bert/bert-base-uncased |
| BERT-large | google-bert/bert-large-uncased |
| RoBERTa-base | FacebookAI/roberta-base |
| RoBERTa-large | FacebookAI/roberta-large |
| ALBERT-base-v2 | albert/albert-base-v2 |
| ALBERT-xlarge-v2 | albert/albert-xlarge-v2 |
Detailed information on the model is in Table 3, which contains the Hugging Face ID of each model.
B.2 Dataset information
We infer using test sets for all benchmarks.
LogiQA
is a logical reasoning dataset sourced from the Chinese Civil Service Examination. It contains 651 test samples, each requiring multi-step logical inference over a given context. Questions cover various reasoning types, including categorical reasoning, conditional reasoning, and disjunctive reasoning.
SciQ
is a science question answering dataset containing 1,000 crowdsourced multiple-choice questions spanning physics, chemistry, and biology. Each question is paired with a short supporting passage (averaging 80 words) that provides the necessary context to answer correctly.
RACE (Reading comprehension from examinations)
consists of English reading comprehension questions collected from Chinese middle and high school exams. We utilize both subsets:
RACE-M: Middle school level with 1,436 test samples and simpler passages (250 words on average)
RACE-H: High school level with 3,498 test samples and more complex texts (350 words on average)
Questions span various types, including detail retrieval, inference, vocabulary, and main idea identification.
B.3 Prompting templates and evaluation protocols
We employ distinct prompting strategies and evaluation protocols tailored to each model architecture—Decoder-only (Causal LM), Encoder-only (Masked LM), and Encoder-Decoder (Seq2Seq)—to ensure fair and optimal performance assessment.
B.3.1 Prompting templates
We evaluate two primary template orderings to test context robustness: CQO (Context Question Options) and QOC (Question Options Context).
Decoder-only models
For standard causal language models (e.g., Llama-3, Qwen, Gemma, Mistral), we use an open-ended completion format that guides the model to predict the next token as the answer label.
CQO template
QOC template
Encoder-only models
For masked language models (e.g., BERT, RoBERTa), we utilize a Cloze-style task where the model must predict the token at the [MASK] position.
CQO template
QOC template
Encoder-Decoder models
For sequence-to-sequence models (e.g., FLAN-T5), the full prompt is provided to the encoder. The decoder is initialized to generate the answer. We use the "Full" variant where the entire content resides in the encoder.
CQO template
QOC template
B.3.2 Evaluation protocols
Decoder-only models (likelihood scoring). Instead of relying on unconstrained text generation, which requires complex parsing and may yield invalid outputs, we evaluate models using a constrained likelihood approach. Given the input prompt, we compute the logits for the next token prediction. We extract the logits corresponding to the valid option tokens (‘A’, ‘B’, ‘C’, ‘D’) and their token variations. We apply a Softmax operation over these four values to obtain a normalized probability distribution:
The model’s prediction is the option with the highest probability. This method allows for a precise measurement of the model’s preference even when differences are subtle.
Encoder-only models (masked prediction). We formulate the task as Masked Language Modeling (MLM). The model processes the bidirectional context and predicts the token at the [MASK] position. Similar to the decoder approach, we extract the probabilities for the tokens corresponding to the options ‘A’, ‘B’, ‘C’, and ‘D’. To handle tokenizer differences, we select the maximum logit across case variations (e.g., ’A’ and ’a’) for each option. These maximum logits are then normalized via Softmax to obtain the final probability distribution.
Encoder-Decoder models. For T5-based models, we employ the same evaluation protocol as decoder-only models.
B.4 Inference details
All experiments are conducted with greedy decoding single runs, maximum 16 output tokens, bfloat16 precision, and NVIDIA A6000 GPUs.
Appendix C Additional result
This section provides detailed per-model and per-dataset results for all experiments discussed in the main paper.
C.1 Full model-dataset results
Table˜4 shows CQO and QOC accuracy for all 21 decoder-only models across 4 datasets. The average gap is +14.7%.
| Model | LogiQA | SciQ | RACE-M | RACE-H | Avg | ||||
| CQO | QOC | CQO | QOC | CQO | QOC | CQO | QOC | Gap | |
| LLaMA Family | |||||||||
| Llama-3.1-8B | 39.8 | 33.8 | 96.3 | 91.8 | 77.6 | 47.1 | 73.9 | 48.2 | +16.7 |
| Llama-3.1-8B-Instruct | 42.1 | 36.6 | 97.8 | 97.1 | 86.1 | 64.9 | 82.1 | 61.0 | +12.1 |
| Llama-3.2-1B | 29.6 | 27.0 | 75.8 | 56.0 | 38.4 | 29.5 | 36.0 | 27.2 | +10.0 |
| Llama-3.2-1B-Instruct | 31.0 | 29.6 | 91.9 | 83.5 | 60.7 | 43.2 | 57.3 | 41.2 | +10.8 |
| Llama-3.2-3B | 27.3 | 23.7 | 93.4 | 84.3 | 64.0 | 39.3 | 60.0 | 37.6 | +15.0 |
| Llama-3.2-3B-Instruct | 34.9 | 32.6 | 96.6 | 94.0 | 80.6 | 54.2 | 75.6 | 52.2 | +13.7 |
| Qwen Family | |||||||||
| Qwen2.5-0.5B | 28.7 | 29.2 | 83.6 | 69.0 | 58.4 | 40.0 | 51.4 | 41.3 | +10.7 |
| Qwen2.5-0.5B-Instruct | 27.0 | 27.5 | 93.6 | 78.9 | 58.6 | 38.6 | 52.8 | 39.5 | +11.9 |
| Qwen2.5-1.5B | 40.4 | 33.9 | 97.7 | 91.9 | 81.5 | 52.4 | 76.3 | 53.4 | +16.1 |
| Qwen2.5-1.5B-Instruct | 42.2 | 34.3 | 97.0 | 90.6 | 80.8 | 47.8 | 75.8 | 48.7 | +18.6 |
| Qwen2.5-3B | 42.5 | 33.0 | 97.7 | 93.2 | 87.3 | 55.9 | 82.4 | 57.0 | +17.7 |
| Qwen2.5-3B-Instruct | 42.9 | 34.4 | 97.8 | 93.8 | 86.8 | 57.6 | 82.2 | 56.8 | +16.8 |
| Qwen2.5-7B | 51.0 | 38.2 | 98.3 | 96.2 | 90.6 | 62.8 | 87.7 | 64.7 | +16.4 |
| Qwen2.5-7B-Instruct | 53.1 | 39.8 | 98.6 | 97.4 | 90.1 | 68.1 | 87.3 | 65.2 | +14.7 |
| Qwen3-0.6B | 37.2 | 26.4 | 87.4 | 69.8 | 51.2 | 33.9 | 46.9 | 32.2 | +15.1 |
| Qwen3-1.7B | 41.0 | 36.7 | 94.7 | 86.3 | 74.9 | 44.4 | 70.9 | 45.6 | +17.1 |
| Qwen3-4B | 52.1 | 39.3 | 98.1 | 95.5 | 86.8 | 55.4 | 80.8 | 55.2 | +18.1 |
| Gemma Family | |||||||||
| gemma-2-2b | 29.0 | 28.4 | 88.1 | 73.3 | 55.4 | 32.1 | 45.5 | 30.6 | +13.4 |
| gemma-2-2b-it | 38.9 | 33.9 | 96.5 | 90.3 | 75.4 | 47.6 | 67.0 | 45.1 | +15.2 |
| gemma-2-9b | 39.2 | 35.8 | 98.0 | 94.9 | 84.8 | 52.4 | 80.2 | 54.1 | +16.3 |
| gemma-2-9b-it | 50.7 | 38.1 | 98.4 | 97.0 | 90.9 | 73.8 | 86.8 | 66.9 | +12.7 |
| Average | 39.1 | 33.0 | 94.1 | 86.9 | 74.3 | 49.6 | 69.5 | 48.8 | +14.7 |
C.2 Base vs instruction-tuned models
Table˜5 compares CQO-QOC gaps between base and instruction-tuned model pairs. Both variants show consistent gaps (Base: 14.70%, Instruct: 14.12%), indicating that instruction tuning does not mitigate order sensitivity.
| Model Pair | Dataset | Base Gap | Inst Gap | Diff |
| LLaMA Family | ||||
| Llama-3.1-8B | LogiQA | +5.99 | +5.53 | -0.46 |
| SciQ | +4.50 | +0.70 | -3.80 | |
| RACE-M | +30.57 | +21.17 | -9.40 | |
| RACE-H | +25.64 | +21.10 | -4.55 | |
| Average | +16.68 | +12.12 | -4.55 | |
| Llama-3.2-1B | LogiQA | +2.61 | +1.38 | -1.23 |
| SciQ | +19.80 | +8.40 | -11.40 | |
| RACE-M | +8.91 | +17.48 | +8.57 | |
| RACE-H | +8.75 | +16.09 | +7.35 | |
| Average | +10.02 | +10.84 | +0.82 | |
| Llama-3.2-3B | LogiQA | +3.69 | +2.30 | -1.38 |
| SciQ | +9.10 | +2.60 | -6.50 | |
| RACE-M | +24.65 | +26.39 | +1.74 | |
| RACE-H | +22.41 | +23.41 | +1.00 | |
| Average | +14.96 | +13.68 | -1.29 | |
| Qwen Family | ||||
| Qwen2.5-0.5B | LogiQA | -0.46 | -0.46 | +0.00 |
| SciQ | +14.60 | +14.70 | +0.10 | |
| RACE-M | +18.31 | +19.99 | +1.67 | |
| RACE-H | +10.15 | +13.32 | +3.17 | |
| Average | +10.65 | +11.89 | +1.24 | |
| Qwen2.5-1.5B | LogiQA | +6.45 | +7.99 | +1.54 |
| SciQ | +5.80 | +6.40 | +0.60 | |
| RACE-M | +29.11 | +32.94 | +3.83 | |
| RACE-H | +22.90 | +27.10 | +4.20 | |
| Average | +16.06 | +18.61 | +2.54 | |
| Qwen2.5-3B | LogiQA | +9.52 | +10.29 | +0.77 |
| SciQ | +4.50 | +4.20 | -0.30 | |
| RACE-M | +31.96 | +29.25 | -2.72 | |
| RACE-H | +25.24 | +25.36 | +0.11 | |
| Average | +17.81 | +17.27 | -0.53 | |
| Qwen2.5-7B | LogiQA | +12.75 | +13.36 | +0.61 |
| SciQ | +2.10 | +1.20 | -0.90 | |
| RACE-M | +27.79 | +22.01 | -5.78 | |
| RACE-H | +23.04 | +22.10 | -0.94 | |
| Average | +16.42 | +14.67 | -1.75 | |
| Gemma Family | ||||
| Gemma-2-2B | LogiQA | +0.61 | +4.92 | +4.30 |
| SciQ | +14.80 | +6.20 | -8.60 | |
| RACE-M | +23.26 | +27.79 | +4.53 | |
| RACE-H | +14.89 | +21.98 | +7.09 | |
| Average | +13.39 | +15.22 | +1.83 | |
| Gemma-2-9B | LogiQA | +3.38 | +12.60 | +9.22 |
| SciQ | +3.10 | +1.40 | -1.70 | |
| RACE-M | +32.45 | +17.06 | -15.39 | |
| RACE-H | +26.13 | +19.93 | -6.20 | |
| Average | +16.26 | +12.75 | -3.52 | |
| Grand Average | – | +14.70 | +14.12 | -0.58 |
C.3 In-context learning results
Table˜6 shows QOC performance with 0, 1, 3, and 5 in-context examples. ICL provides marginal improvement (+3.1% from 0-shot to 5-shot) but cannot close the gap to CQO (69.26%).
| Model | 0-shot | 1-shot | 3-shot | 5-shot | |
| LLaMA Family | |||||
| Llama-3.1-8B | 55.22 | 60.35 | 62.29 | 62.78 | +7.56 |
| Llama-3.1-8B-Instruct | 64.90 | 64.30 | 65.67 | 64.80 | -0.10 |
| Llama-3.2-1B | 34.93 | 28.05 | 31.93 | 32.64 | -2.30 |
| Llama-3.2-1B-Instruct | 49.39 | 46.59 | 50.68 | 51.54 | +2.15 |
| Llama-3.2-3B | 46.22 | 50.18 | 53.24 | 53.27 | +7.04 |
| Llama-3.2-3B-Instruct | 58.24 | 57.23 | 58.78 | 59.60 | +1.35 |
| Qwen Family | |||||
| Qwen2.5-0.5B | 44.87 | 44.29 | 45.06 | 45.37 | +0.50 |
| Qwen2.5-0.5B-Instruct | 46.13 | 41.46 | 40.30 | 42.45 | -3.68 |
| Qwen2.5-1.5B | 57.90 | 57.60 | 58.45 | 58.55 | +0.65 |
| Qwen2.5-1.5B-Instruct | 55.36 | 54.70 | 57.26 | 58.14 | +2.78 |
| Qwen2.5-3B | 59.75 | 59.49 | 61.26 | 61.06 | +1.31 |
| Qwen2.5-3B-Instruct | 60.49 | 59.22 | 61.36 | 61.92 | +1.42 |
| Qwen2.5-7B | 65.49 | 65.95 | 66.52 | 65.82 | +0.33 |
| Qwen2.5-7B-Instruct | 67.62 | 67.19 | 67.71 | 68.50 | +0.88 |
| Qwen3-0.6B | 40.60 | 44.08 | 50.17 | 51.65 | +11.05 |
| Qwen3-1.7B | 53.25 | 55.13 | 58.51 | 58.79 | +5.54 |
| Qwen3-4B | 61.36 | 64.86 | 67.25 | 67.15 | +5.79 |
| Gemma Family | |||||
| Gemma-2-2B | 41.10 | 47.84 | 49.94 | 49.89 | +8.80 |
| Gemma-2-2B-Instruct | 54.23 | 59.89 | 61.01 | 61.26 | +7.03 |
| Gemma-2-9B | 59.29 | 62.33 | 64.46 | 65.28 | +5.99 |
| Gemma-2-9B-Instruct | 68.95 | 67.72 | 70.09 | 69.75 | +0.79 |
| Average | 54.54 | 55.16 | 57.24 | 57.63 | +3.09 |
C.4 Option recall accuracy
Table˜7 shows option recall accuracy by dataset. High recall rates in both CQO (93.5%) and QOC (94.7%) confirm that the QOC performance drop is not due to memory failure.
| LogiQA | SciQ | RACE-M | RACE-H | Avg | ||||||
| Model | CQO | QOC | CQO | QOC | CQO | QOC | CQO | QOC | CQO | QOC |
| Llama-3.1-8B | 93.3 | 94.5 | 99.8 | 99.8 | 97.3 | 97.2 | 96.9 | 96.9 | 96.8 | 97.1 |
| Llama-3.1-8B-Instruct | 96.7 | 96.8 | 99.8 | 99.8 | 97.0 | 98.3 | 96.5 | 98.5 | 97.5 | 98.3 |
| Llama-3.2-1B | 89.9 | 85.3 | 97.8 | 97.6 | 87.2 | 88.1 | 87.4 | 88.2 | 90.6 | 89.8 |
| Llama-3.2-1B-Instruct | 89.7 | 91.2 | 98.3 | 99.1 | 92.0 | 97.6 | 91.9 | 97.7 | 93.0 | 96.4 |
| Llama-3.2-3B | 94.0 | 92.9 | 99.6 | 98.0 | 95.4 | 94.0 | 94.5 | 92.8 | 95.9 | 94.4 |
| Llama-3.2-3B-Instruct | 93.1 | 93.7 | 95.2 | 99.1 | 91.2 | 97.9 | 88.4 | 97.8 | 92.0 | 97.1 |
| Qwen2.5-0.5B | 89.6 | 90.0 | 99.2 | 98.4 | 96.5 | 96.3 | 96.8 | 96.1 | 95.5 | 95.2 |
| Qwen2.5-0.5B-Instruct | 85.9 | 86.1 | 94.9 | 96.0 | 92.5 | 95.7 | 92.9 | 95.1 | 91.6 | 93.2 |
| Qwen2.5-1.5B | 88.7 | 87.7 | 93.8 | 96.2 | 93.4 | 95.5 | 94.9 | 95.2 | 92.7 | 93.6 |
| Qwen2.5-1.5B-Instruct | 86.9 | 85.6 | 95.0 | 97.0 | 91.4 | 94.4 | 92.3 | 95.0 | 91.4 | 93.0 |
| Qwen2.5-3B | 88.9 | 89.0 | 99.2 | 99.2 | 95.9 | 96.8 | 95.8 | 96.5 | 95.0 | 95.4 |
| Qwen2.5-3B-Instruct | 88.7 | 89.3 | 99.1 | 99.2 | 96.5 | 97.4 | 96.2 | 96.9 | 95.1 | 95.7 |
| Qwen2.5-7B | 89.7 | 89.6 | 99.5 | 99.4 | 96.4 | 96.6 | 96.2 | 96.4 | 95.5 | 95.5 |
| Qwen2.5-7B-Instruct | 86.9 | 87.3 | 99.2 | 99.4 | 96.2 | 96.8 | 95.7 | 96.1 | 94.5 | 94.9 |
| Qwen3-0.6B | 87.7 | 89.6 | 96.5 | 97.0 | 77.3 | 91.5 | 84.2 | 93.3 | 86.4 | 92.8 |
| Qwen3-1.7B | 91.0 | 92.7 | 99.3 | 99.6 | 97.0 | 97.9 | 97.3 | 98.1 | 96.1 | 97.1 |
| Qwen3-4B | 88.9 | 89.2 | 99.7 | 99.6 | 97.5 | 97.7 | 97.2 | 97.4 | 95.8 | 96.0 |
| gemma-2-2b-it | 84.0 | 84.3 | 93.5 | 89.6 | 76.7 | 84.3 | 85.0 | 89.5 | 84.8 | 86.9 |
| gemma-2-9b-it | 94.8 | 95.1 | 98.8 | 99.2 | 97.1 | 98.0 | 96.4 | 97.8 | 96.8 | 97.5 |
| Average | 89.9 | 90.0 | 97.8 | 98.1 | 92.9 | 95.4 | 93.5 | 95.5 | 93.5 | 94.7 |
C.5 Encoder and encoder-decoder model results
Table˜8 shows results for encoder-only models and encoder-decoder models.
| Model | Dataset | CQO | QOC | Gap |
| Encoder Only Models | ||||
| BERT-Base | LogiQA | 20.89 | 19.20 | +1.69 |
| SciQ | 24.60 | 25.60 | -1.00 | |
| RACE-M | 21.10 | 21.38 | -0.28 | |
| RACE-H | 18.95 | 17.92 | +1.03 | |
| Average | 21.39 | 21.03 | +0.36 | |
| BERT-Large | LogiQA | 20.74 | 19.82 | +0.92 |
| SciQ | 23.90 | 26.50 | -2.60 | |
| RACE-M | 23.68 | 21.73 | +1.95 | |
| RACE-H | 21.04 | 18.78 | +2.26 | |
| Average | 22.34 | 21.71 | +0.63 | |
| RoBERTa-Base | LogiQA | 20.74 | 20.74 | +0.00 |
| SciQ | 28.90 | 27.80 | +1.10 | |
| RACE-M | 21.17 | 20.89 | +0.28 | |
| RACE-H | 16.67 | 16.41 | +0.26 | |
| Average | 21.87 | 21.46 | +0.41 | |
| RoBERTa-Large | LogiQA | 21.97 | 25.19 | -3.23 |
| SciQ | 28.60 | 28.90 | -0.30 | |
| RACE-M | 21.73 | 23.47 | -1.74 | |
| RACE-H | 16.87 | 17.50 | -0.63 | |
| Average | 22.29 | 23.76 | -1.47 | |
| ALBERT-Base | LogiQA | 20.28 | 20.12 | +0.15 |
| SciQ | 25.60 | 25.90 | -0.30 | |
| RACE-M | 20.68 | 20.82 | -0.14 | |
| RACE-H | 17.21 | 17.21 | +0.00 | |
| Average | 20.94 | 21.01 | -0.07 | |
| ALBERT-XLarge | LogiQA | 19.97 | 20.12 | -0.15 |
| SciQ | 26.30 | 26.10 | +0.20 | |
| RACE-M | 20.82 | 20.61 | +0.21 | |
| RACE-H | 17.30 | 17.27 | +0.03 | |
| Average | 21.10 | 21.03 | +0.07 | |
| Encoder Avg | – | – | – | -0.02 |
| Model | Dataset | CQO | QOC | Gap |
| Encoder-Decoder Models | ||||
| Flan-T5-Small | LogiQA | 27.80 | 29.95 | -2.15 |
| SciQ | 88.70 | 84.20 | +4.50 | |
| RACE-M | 51.74 | 44.64 | +7.10 | |
| RACE-H | 42.71 | 38.02 | +4.69 | |
| Average | 52.74 | 49.20 | +3.54 | |
| Flan-T5-Base | LogiQA | 32.10 | 30.41 | +1.69 |
| SciQ | 94.60 | 93.60 | +1.00 | |
| RACE-M | 76.25 | 72.91 | +3.34 | |
| RACE-H | 66.52 | 63.15 | +3.37 | |
| Average | 67.37 | 65.02 | +2.35 | |
| Flan-T5-Large | LogiQA | 36.41 | 34.10 | +2.30 |
| SciQ | 96.40 | 96.00 | +0.40 | |
| RACE-M | 85.31 | 84.19 | +1.11 | |
| RACE-H | 80.45 | 76.56 | +3.89 | |
| Average | 74.64 | 72.71 | +1.93 | |
| Flan-T5-XL | LogiQA | 35.33 | 33.33 | +2.00 |
| SciQ | 97.40 | 97.10 | +0.30 | |
| RACE-M | 90.18 | 88.86 | +1.32 | |
| RACE-H | 86.68 | 84.73 | +1.94 | |
| Average | 77.40 | 76.01 | +1.39 | |
| Enc-Dec Avg | – | – | – | +2.30 |
C.6 Gradient attribution
Figure˜7 shows context attribution ratios by dataset. Across all datasets, CQO receives significantly more gradient flow from context tokens than QOC (average ratio: 2.38). We also show details in Table˜9.
| Model | Dataset | CQO | QOC | Ratio |
| LLaMA Family | ||||
| Llama-3.1-8B | LogiQA | 0.806 | 0.257 | 3.133 |
| SciQ | 0.769 | 0.269 | 2.859 | |
| RACE-M | 0.859 | 0.368 | 2.335 | |
| RACE-H | 0.881 | 0.400 | 2.205 | |
| Average | 0.829 | 0.323 | 2.562 | |
| Llama-3.1-8B-Instruct | LogiQA | 0.835 | 0.248 | 3.364 |
| SciQ | 0.767 | 0.294 | 2.612 | |
| RACE-M | 0.870 | 0.403 | 2.157 | |
| RACE-H | 0.891 | 0.432 | 2.060 | |
| Average | 0.841 | 0.344 | 2.441 | |
| Llama-3.2-1B | LogiQA | 0.797 | 0.255 | 3.129 |
| SciQ | 0.814 | 0.220 | 3.697 | |
| RACE-M | 0.868 | 0.386 | 2.251 | |
| RACE-H | 0.881 | 0.414 | 2.128 | |
| Average | 0.840 | 0.319 | 2.637 | |
| Llama-3.2-1B-Instruct | LogiQA | 0.841 | 0.293 | 2.869 |
| SciQ | 0.808 | 0.299 | 2.700 | |
| RACE-M | 0.889 | 0.413 | 2.154 | |
| RACE-H | 0.906 | 0.424 | 2.137 | |
| Average | 0.861 | 0.357 | 2.410 | |
| Llama-3.2-3B | LogiQA | 0.836 | 0.275 | 3.037 |
| SciQ | 0.804 | 0.267 | 3.007 | |
| RACE-M | 0.868 | 0.405 | 2.142 | |
| RACE-H | 0.893 | 0.429 | 2.080 | |
| Average | 0.850 | 0.344 | 2.470 | |
| Llama-3.2-3B-Instruct | LogiQA | 0.877 | 0.298 | 2.942 |
| SciQ | 0.815 | 0.319 | 2.556 | |
| RACE-M | 0.901 | 0.460 | 1.957 | |
| RACE-H | 0.917 | 0.485 | 1.892 | |
| Average | 0.877 | 0.390 | 2.247 | |
| Qwen Family | ||||
| Qwen2.5-0.5B | LogiQA | 0.855 | 0.306 | 2.798 |
| SciQ | 0.835 | 0.282 | 2.958 | |
| RACE-M | 0.881 | 0.375 | 2.350 | |
| RACE-H | 0.890 | 0.407 | 2.187 | |
| Average | 0.865 | 0.342 | 2.526 | |
| Qwen2.5-0.5B-Instruct | LogiQA | 0.781 | 0.263 | 2.973 |
| SciQ | 0.816 | 0.250 | 3.260 | |
| RACE-M | 0.843 | 0.399 | 2.111 | |
| RACE-H | 0.850 | 0.428 | 1.987 | |
| Average | 0.823 | 0.335 | 2.455 | |
| Qwen2.5-1.5B | LogiQA | 0.876 | 0.250 | 3.505 |
| SciQ | 0.823 | 0.219 | 3.752 | |
| RACE-M | 0.884 | 0.312 | 2.834 | |
| RACE-H | 0.892 | 0.325 | 2.745 | |
| Average | 0.869 | 0.277 | 3.142 | |
| Qwen2.5-1.5B-Instruct | LogiQA | 0.825 | 0.210 | 3.926 |
| SciQ | 0.775 | 0.236 | 3.285 | |
| RACE-M | 0.805 | 0.355 | 2.269 | |
| RACE-H | 0.820 | 0.373 | 2.197 | |
| Average | 0.806 | 0.293 | 2.747 | |
| Model | Dataset | CQO | QOC | Ratio |
| Qwen Family (cont.) | ||||
| Qwen2.5-3B | LogiQA | 0.871 | 0.224 | 3.886 |
| SciQ | 0.797 | 0.276 | 2.889 | |
| RACE-M | 0.849 | 0.351 | 2.421 | |
| RACE-H | 0.872 | 0.342 | 2.548 | |
| Average | 0.848 | 0.298 | 2.841 | |
| Qwen2.5-3B-Instruct | LogiQA | 0.881 | 0.194 | 4.532 |
| SciQ | 0.832 | 0.281 | 2.965 | |
| RACE-M | 0.873 | 0.374 | 2.333 | |
| RACE-H | 0.890 | 0.387 | 2.298 | |
| Average | 0.869 | 0.309 | 2.811 | |
| Qwen2.5-7B | LogiQA | 0.867 | 0.338 | 2.564 |
| SciQ | 0.868 | 0.368 | 2.360 | |
| RACE-M | 0.890 | 0.479 | 1.859 | |
| RACE-H | 0.907 | 0.500 | 1.814 | |
| Average | 0.883 | 0.421 | 2.097 | |
| Qwen2.5-7B-Instruct | LogiQA | 0.857 | 0.256 | 3.350 |
| SciQ | 0.845 | 0.366 | 2.310 | |
| RACE-M | 0.870 | 0.473 | 1.840 | |
| RACE-H | 0.875 | 0.491 | 1.781 | |
| Average | 0.862 | 0.396 | 2.174 | |
| Qwen3-0.6B | LogiQA | 0.860 | 0.189 | 4.553 |
| SciQ | 0.832 | 0.219 | 3.806 | |
| RACE-M | 0.852 | 0.316 | 2.695 | |
| RACE-H | 0.869 | 0.341 | 2.547 | |
| Average | 0.853 | 0.266 | 3.206 | |
| Qwen3-1.7B | LogiQA | 0.849 | 0.197 | 4.301 |
| SciQ | 0.791 | 0.256 | 3.083 | |
| RACE-M | 0.859 | 0.357 | 2.408 | |
| RACE-H | 0.871 | 0.364 | 2.390 | |
| Average | 0.842 | 0.294 | 2.868 | |
| Qwen3-4B | LogiQA | 0.928 | 0.179 | 5.178 |
| SciQ | 0.846 | 0.318 | 2.658 | |
| RACE-M | 0.897 | 0.446 | 2.010 | |
| RACE-H | 0.902 | 0.430 | 2.097 | |
| Average | 0.893 | 0.343 | 2.601 | |
| Gemma Family | ||||
| Gemma-2-2B | LogiQA | 0.513 | 0.276 | 1.861 |
| SciQ | 0.480 | 0.233 | 2.061 | |
| RACE-M | 0.589 | 0.361 | 1.633 | |
| RACE-H | 0.604 | 0.396 | 1.524 | |
| Average | 0.547 | 0.316 | 1.727 | |
| Gemma-2-2B-Instruct | LogiQA | 0.445 | 0.199 | 2.233 |
| SciQ | 0.462 | 0.294 | 1.570 | |
| RACE-M | 0.582 | 0.406 | 1.433 | |
| RACE-H | 0.593 | 0.441 | 1.345 | |
| Average | 0.520 | 0.335 | 1.553 | |
| Gemma-2-9B | LogiQA | 0.519 | 0.247 | 2.095 |
| SciQ | 0.519 | 0.267 | 1.947 | |
| RACE-M | 0.617 | 0.373 | 1.655 | |
| RACE-H | 0.633 | 0.396 | 1.600 | |
| Average | 0.572 | 0.321 | 1.784 | |
| Gemma-2-9B-Instruct | LogiQA | 0.524 | 0.262 | 2.000 |
| SciQ | 0.558 | 0.332 | 1.681 | |
| RACE-M | 0.654 | 0.499 | 1.310 | |
| RACE-H | 0.655 | 0.553 | 1.186 | |
| Average | 0.598 | 0.411 | 1.453 | |
| Grand Average | – | 0.797 | 0.335 | 2.380 |
C.7 Intervention results
Attention pruning (CQO Degradation).
Table˜10 shows the effect of blocking option-to-context attention in CQO prompts. The average accuracy drop is -26.8%, confirming this attention pathway is essential.
| Model | LogiQA | SciQ | RACE-M | RACE-H | Avg | ||||||||
| CQO | Pruned | CQO | Pruned | CQO | Pruned | CQO | Pruned | ||||||
| LLaMA Family | |||||||||||||
| Llama-3.1-8B | 39.78 | 30.88 | -8.91 | 96.30 | 89.90 | -6.40 | 77.65 | 51.95 | -25.70 | 73.87 | 49.60 | -24.27 | -16.32 |
| Llama-3.1-8B-Instruct | 42.09 | 35.79 | -6.30 | 97.80 | 93.70 | -4.10 | 86.07 | 58.98 | -27.09 | 82.13 | 58.75 | -23.38 | -15.22 |
| Llama-3.2-1B | 29.65 | 29.34 | -0.31 | 75.80 | 61.50 | -14.30 | 38.37 | 32.03 | -6.34 | 35.99 | 31.30 | -4.69 | -6.41 |
| Llama-3.2-1B-Instruct | 31.03 | 31.18 | +0.15 | 91.90 | 83.30 | -8.60 | 60.72 | 42.34 | -18.38 | 57.26 | 42.65 | -14.61 | -10.36 |
| Llama-3.2-3B | 27.34 | 30.41 | +3.07 | 93.40 | 76.70 | -16.70 | 64.00 | 42.69 | -21.31 | 60.01 | 41.80 | -18.21 | -13.29 |
| Llama-3.2-3B-Instruct | 34.87 | 33.64 | -1.23 | 96.60 | 90.00 | -6.60 | 80.57 | 53.83 | -26.74 | 75.64 | 53.72 | -21.93 | -14.12 |
| Qwen Family | |||||||||||||
| Qwen2.5-0.5B | 28.73 | 27.34 | -1.38 | 83.60 | 25.60 | -58.00 | 58.36 | 26.60 | -31.75 | 51.40 | 25.96 | -25.44 | -29.15 |
| Qwen2.5-0.5B-Instruct | 27.04 | 26.27 | -0.77 | 93.60 | 26.30 | -67.30 | 58.64 | 25.77 | -32.87 | 52.80 | 25.07 | -27.73 | -32.17 |
| Qwen2.5-1.5B | 40.40 | 27.80 | -12.60 | 97.70 | 29.70 | -68.00 | 81.48 | 26.74 | -54.74 | 76.27 | 25.53 | -50.74 | -46.52 |
| Qwen2.5-1.5B-Instruct | 42.24 | 25.81 | -16.44 | 97.00 | 31.50 | -65.50 | 80.78 | 27.09 | -53.69 | 75.84 | 26.44 | -49.40 | -46.26 |
| Qwen2.5-3B | 42.55 | 27.34 | -15.21 | 97.80 | 24.50 | -73.30 | 87.47 | 26.32 | -61.14 | 82.42 | 24.87 | -57.55 | -51.80 |
| Qwen2.5-3B-Instruct | 44.09 | 28.57 | -15.51 | 97.80 | 25.60 | -72.20 | 86.91 | 26.67 | -60.24 | 82.28 | 25.90 | -56.38 | -51.08 |
| Qwen2.5-7B | 51.00 | 31.03 | -19.97 | 98.30 | 79.90 | -18.40 | 90.60 | 48.89 | -41.71 | 87.74 | 45.03 | -42.71 | -30.70 |
| Qwen2.5-7B-Instruct | 53.15 | 35.33 | -17.82 | 98.60 | 81.30 | -17.30 | 90.11 | 46.87 | -43.25 | 87.31 | 46.20 | -41.11 | -29.87 |
| Qwen3-0.6B | 37.17 | 21.97 | -15.21 | 87.40 | 27.30 | -60.10 | 51.18 | 21.59 | -29.60 | 46.88 | 21.47 | -25.41 | -32.58 |
| Qwen3-1.7B | 41.01 | 26.57 | -14.44 | 94.87 | 30.40 | -64.47 | 74.86 | 26.32 | -48.54 | 70.87 | 27.90 | -42.97 | -42.60 |
| Qwen3-4B | 52.07 | 26.42 | -25.65 | 98.10 | 43.10 | -55.00 | 86.84 | 26.39 | -60.45 | 80.85 | 28.67 | -52.17 | -48.32 |
| Gemma Family | |||||||||||||
| Gemma-2-2B | 29.03 | 27.65 | -1.38 | 88.10 | 80.60 | -7.50 | 55.36 | 39.55 | -15.81 | 45.45 | 40.42 | -5.03 | -7.43 |
| Gemma-2-2B-Instruct | 38.86 | 37.79 | -1.08 | 96.50 | 89.90 | -6.60 | 75.42 | 52.37 | -23.05 | 67.04 | 53.06 | -13.98 | -11.18 |
| Gemma-2-9B | 39.17 | 35.64 | -3.53 | 98.00 | 94.40 | -3.60 | 84.82 | 58.70 | -26.11 | 80.25 | 57.86 | -22.38 | -13.91 |
| Gemma-2-9B-Instruct | 50.69 | 43.32 | -7.37 | 98.40 | 94.90 | -3.50 | 90.88 | 68.25 | -22.63 | 86.82 | 66.92 | -19.90 | -13.35 |
| Average | 39.14 | 30.48 | -8.66 | 94.17 | 60.96 | -33.21 | 74.34 | 39.52 | -34.82 | 69.48 | 39.01 | -30.48 | -26.79 |
Activation patching (QOC Improvement).
Table˜11 shows the effect of replacing QOC option hidden states with CQO representations. Average improvement: +6.0%.
| Model | LogiQA | SciQ | RACE-M | RACE-H | Avg | ||||
| QOC | Patched | QOC | Patched | QOC | Patched | QOC | Patched | ||
| LLaMA Family | |||||||||
| Llama-3.1-8B | 33.79 | 30.57 | 91.80 | 87.20 | 47.08 | 52.58 | 48.23 | 48.77 | -0.45 |
| Llama-3.1-8B-Instruct | 36.56 | 28.42 | 97.10 | 94.70 | 64.90 | 61.98 | 61.03 | 51.97 | -5.63 |
| Llama-3.2-1B | 27.04 | 26.73 | 56.00 | 57.00 | 29.46 | 29.53 | 27.24 | 28.70 | +0.56 |
| Llama-3.2-1B-Instruct | 29.65 | 30.11 | 83.50 | 82.40 | 43.25 | 47.08 | 41.17 | 43.77 | +1.45 |
| Llama-3.2-3B | 23.66 | 30.72 | 84.30 | 73.20 | 39.35 | 41.09 | 37.59 | 36.45 | -0.86 |
| Llama-3.2-3B-Instruct | 32.57 | 31.80 | 94.00 | 93.40 | 54.18 | 55.78 | 52.23 | 47.97 | -1.01 |
| Qwen Family | |||||||||
| Qwen2.5-0.5B | 29.19 | 26.42 | 69.00 | 76.60 | 40.04 | 41.71 | 41.25 | 37.14 | +0.60 |
| Qwen2.5-0.5B-Instruct | 27.50 | 23.96 | 78.90 | 78.70 | 38.65 | 36.98 | 39.48 | 32.76 | -3.03 |
| Qwen2.5-1.5B | 33.95 | 37.02 | 91.90 | 95.00 | 52.37 | 71.24 | 53.37 | 64.69 | +9.09 |
| Qwen2.5-1.5B-Instruct | 34.25 | 37.48 | 90.60 | 92.10 | 47.84 | 67.34 | 48.74 | 61.61 | +9.27 |
| Qwen2.5-3B | 33.03 | 34.72 | 93.30 | 93.30 | 55.50 | 70.96 | 57.18 | 66.98 | +6.74 |
| Qwen2.5-3B-Instruct | 33.79 | 31.34 | 93.60 | 96.20 | 57.66 | 72.70 | 56.92 | 66.21 | +6.12 |
| Qwen2.5-7B | 38.25 | 46.08 | 96.20 | 98.30 | 62.81 | 88.72 | 64.69 | 83.28 | +13.61 |
| Qwen2.5-7B-Instruct | 39.78 | 52.38 | 97.40 | 98.10 | 68.11 | 87.67 | 65.21 | 83.13 | +12.70 |
| Qwen3-0.6B | 26.42 | 31.95 | 69.80 | 83.20 | 33.91 | 56.27 | 32.25 | 50.71 | +14.94 |
| Qwen3-1.7B | 36.71 | 38.86 | 86.30 | 92.00 | 44.43 | 71.80 | 45.57 | 68.21 | +14.46 |
| Qwen3-4B | 39.32 | 50.23 | 95.50 | 97.90 | 55.43 | 84.96 | 55.17 | 78.73 | +16.60 |
| Gemma Family | |||||||||
| Gemma-2-2B | 28.42 | 28.26 | 73.30 | 89.10 | 32.10 | 53.06 | 30.56 | 45.28 | +12.83 |
| Gemma-2-2B-Instruct | 33.95 | 38.56 | 90.30 | 94.50 | 47.63 | 72.70 | 45.05 | 67.64 | +14.12 |
| Gemma-2-9B | 35.79 | 32.72 | 94.90 | 71.80 | 52.37 | 62.12 | 54.12 | 59.23 | -2.83 |
| Gemma-2-9B-Instruct | 38.10 | 50.38 | 97.00 | 95.80 | 73.82 | 80.43 | 66.90 | 72.44 | +5.81 |
| Average | 33.42 | 34.76 | 89.62 | 90.96 | 49.72 | 65.69 | 48.78 | 61.48 | +5.96 |
Option repetition (QOCO).
Table˜12 shows the effect of repeating options after context (Q-O-C-O format). This simple modification allows the repeated options to attend to context, improving QOC accuracy by an average of +8.2%.
| Model | LogiQA | SciQ | RACE-M | RACE-H | Avg | ||||||||
| QOC | QOCO | QOC | QOCO | QOC | QOCO | QOC | QOCO | ||||||
| LLaMA Family | |||||||||||||
| Llama-3.1-8B | 33.79 | 38.40 | +4.61 | 91.80 | 86.90 | -4.90 | 47.08 | 59.19 | +12.11 | 48.23 | 58.98 | +10.75 | +5.64 |
| Llama-3.1-8B-Instruct | 36.56 | 39.48 | +2.92 | 97.10 | 97.40 | +0.30 | 64.90 | 78.55 | +13.65 | 61.03 | 73.79 | +12.76 | +7.40 |
| Llama-3.2-1B | 27.04 | 27.19 | +0.15 | 56.00 | 68.50 | +12.50 | 29.46 | 34.26 | +4.80 | 27.24 | 32.08 | +4.84 | +5.57 |
| Llama-3.2-1B-Instruct | 29.65 | 29.03 | -0.62 | 83.50 | 94.20 | +10.70 | 43.25 | 55.85 | +12.60 | 41.17 | 51.43 | +10.26 | +8.23 |
| Llama-3.2-3B | 23.66 | 30.57 | +6.91 | 84.30 | 92.40 | +8.10 | 39.35 | 53.76 | +14.41 | 37.59 | 47.66 | +10.07 | +9.87 |
| Llama-3.2-3B-Instruct | 32.57 | 35.48 | +2.91 | 94.00 | 96.50 | +2.50 | 54.18 | 72.14 | +17.96 | 52.23 | 66.47 | +14.24 | +9.40 |
| Qwen Family | |||||||||||||
| Qwen2.5-0.5B | 29.19 | 27.50 | -1.69 | 69.00 | 85.10 | +16.10 | 40.04 | 43.31 | +3.27 | 41.25 | 38.74 | -2.51 | +3.79 |
| Qwen2.5-0.5B-Instruct | 27.50 | 24.27 | -3.23 | 78.90 | 88.70 | +9.80 | 38.65 | 48.40 | +9.75 | 39.48 | 44.57 | +5.09 | +5.35 |
| Qwen2.5-1.5B | 33.95 | 33.95 | 0.00 | 91.90 | 95.60 | +3.70 | 52.37 | 63.58 | +11.21 | 53.37 | 56.43 | +3.06 | +4.49 |
| Qwen2.5-1.5B-Instruct | 34.25 | 35.02 | +0.77 | 90.60 | 91.40 | +0.80 | 47.84 | 60.38 | +12.54 | 48.74 | 59.75 | +11.01 | +6.28 |
| Qwen2.5-3B | 33.03 | 35.94 | +2.91 | 93.30 | 97.10 | +3.80 | 55.50 | 71.31 | +15.81 | 57.18 | 66.18 | +9.00 | +7.88 |
| Qwen2.5-3B-Instruct | 33.79 | 38.56 | +4.77 | 93.60 | 96.80 | +3.20 | 57.66 | 73.75 | +16.09 | 56.92 | 69.01 | +12.09 | +9.04 |
| Qwen2.5-7B | 38.25 | 42.70 | +4.45 | 96.20 | 98.00 | +1.80 | 62.81 | 79.18 | +16.37 | 64.69 | 74.76 | +10.07 | +8.17 |
| Qwen2.5-7B-Instruct | 39.78 | 45.78 | +6.00 | 97.40 | 97.50 | +0.10 | 68.11 | 80.99 | +12.88 | 65.21 | 76.90 | +11.69 | +7.66 |
| Qwen3-0.6B | 26.42 | 34.56 | +8.14 | 69.80 | 80.10 | +10.30 | 33.91 | 49.86 | +15.95 | 32.25 | 45.20 | +12.95 | +11.83 |
| Qwen3-1.7B | 36.71 | 36.10 | -0.61 | 86.30 | 82.10 | -4.20 | 44.43 | 55.99 | +11.56 | 45.57 | 56.98 | +11.41 | +4.54 |
| Qwen3-4B | 39.32 | 48.54 | +9.22 | 95.50 | 98.20 | +2.70 | 55.43 | 76.74 | +21.31 | 55.17 | 72.41 | +17.24 | +12.61 |
| Gemma Family | |||||||||||||
| Gemma-2-2B | 28.42 | 26.11 | -2.31 | 73.30 | 91.60 | +18.30 | 32.10 | 48.05 | +15.95 | 30.56 | 45.25 | +14.69 | +11.65 |
| Gemma-2-2B-Instruct | 33.95 | 41.01 | +7.06 | 90.30 | 94.90 | +4.60 | 47.63 | 71.80 | +24.17 | 45.05 | 65.84 | +20.79 | +14.15 |
| Gemma-2-9B | 35.79 | 38.25 | +2.46 | 94.90 | 97.60 | +2.70 | 52.37 | 71.80 | +19.43 | 54.12 | 69.07 | +14.95 | +9.88 |
| Gemma-2-9B-Instruct | 38.10 | 48.08 | +9.98 | 97.00 | 97.30 | +0.30 | 73.82 | 85.65 | +11.83 | 66.90 | 81.25 | +14.35 | +9.11 |
| Average | 32.93 | 36.02 | +3.08 | 86.89 | 91.80 | +4.91 | 49.56 | 63.54 | +13.98 | 48.75 | 59.65 | +10.89 | +8.21 |